OpenStudy (ineedtheanswersquick):

Solve each equation. #2

OpenStudy (ineedtheanswersquick):

−6 + x/4 = −5

OpenStudy (ineedtheanswersquick):

@AlexandervonHumboldt2

OpenStudy (alexandervonhumboldt2):

add 6 to both sides: -6+6+x/4=-5+6 x/4=1 x=1*4=?